OpenStudy (anonymous):

Which expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = 14x -15 2 (7x +10 ) 15x-14 2(7x-9) Choose exactly 2 answers that are correct please help! will medal and fan :)

Nnesha (nnesha):

okay again distribute parentheses by 6 and then combine like terms :)

OpenStudy (anonymous):

i got 14x -18 and i know d isnt an option :)

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

there are two choices with parentheses so distribute that

Nnesha (nnesha):

\[2(7x +10 ) = ??\] \[2(7x -9 ) = ??\]

OpenStudy (anonymous):

14x +20 and 14x - 18

Nnesha (nnesha):

so which one is your answer :) ??

OpenStudy (anonymous):

14x - 18?

OpenStudy (anonymous):

its 2 answers that equal to no solution

OpenStudy (anonymous):

For it to have no solution, the answer would have to be something like 2=3, where 6(4x-3) = 200x-15 (just an random example) would simply to 2=3

OpenStudy (anonymous):

My example doesn't simplify to 2=3, I just made up a random equation.

OpenStudy (anonymous):

Yeah ! but there's 2 answers to this question and all 3 of them seem like they aren't equal (Except for d)

OpenStudy (anonymous):

Lets start with the right side being 2(7x+10)

OpenStudy (anonymous):

6(4x-3) - 10x = 2(7x+10)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Balvasor 14x - 18? \(\color{blue}{\text{End of Quote}}\) yes this is right

OpenStudy (anonymous):

14x - 18 = 14x + 20

OpenStudy (anonymous):

thanks! and im guessing the second answer is 15x - 14?

Nnesha (nnesha):

how it can be 15x ?

OpenStudy (anonymous):

ok 14x- 15 ? haha

Nnesha (nnesha):

you got 14x - 18 how it can be 15 :/ hmmm

OpenStudy (anonymous):

im supposed to find one that equals no solution

OpenStudy (anonymous):

i already got b like you said but there's supposed to be 2 answers

Nnesha (nnesha):

ohh wait you need the equation which would give you no solution right

Nnesha (nnesha):

sorry i didn't read that part

Nnesha (nnesha):

olay so you know that 6 (4x - 3) - 10x = 14x - 18 so now for D you have 14x -18 = 14x -18 if you get 0 = 0 then infinity many solution if both side are not e equal then no solution

Nnesha (nnesha):

14x + 20 = 14x - 18 solve for x

OpenStudy (anonymous):

no solution?

Nnesha (nnesha):

yep right

Nnesha (nnesha):

14x - 18 = 14x -15 ???

OpenStudy (anonymous):

no solution

Nnesha (nnesha):

Yep :)

OpenStudy (anonymous):

thank you so much!!

Nnesha (nnesha):

np :)